C ^{A}Modules

by

Alvin Zhang

A thesis

presented to the University of Waterloo

in fulfillment of the thesis requirement for the degree of

Master of Mathematics

in Computer Science

Waterloo, Ontario, Canada, 2026

© Alvin Zhang 2026

Author’s Declaration

I hereby declare that I am the sole author of this thesis. This is a true copy of the thesis, including any required final revisions, as accepted by my examiners. I understand that my thesis may be made electronically available to the public.

Abstract

This is the abstract.

Acknowledgements

I would like to thank all the little people who made this thesis possible.

Table of Contents

Author’s Declaration ii Abstract iii Acknowledgements iv List of Figures vi List of Tables vii 1 Introduction 1 2 Background 2 3 Modules 3 4 Conclusion 4 References 5

List of Figures

vi

List of Tables

vii

Chapter 1

Introduction Chapter 2 Background

2

Chapter 3 Modules

3

Chapter 4 Conclusion

4

References

5